# DS-VTON: High-Quality Virtual Try-on via Disentangled Dual-Scale Generation

### v1

@@ 3.4 High-resolution stage
While the low-resolution stage focuses on body structure by ignoring fine details, the high-resolution stage is designed to refine garment textures. Instead of directly reconstructing the full high-resolution image, this stage aims to reconstruct the residual between the high- and low-resolution outputs. This strategy effectively disentangles structural alignment from detailed appearance modeling.
3.4.1 Reformulating the denoising process with residual learning
@@
Equations (2) and (3) define the forward and reverse diffusion processes, respectively. In these equations, $\bm{\epsilon},\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, and $\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)$ denotes the noise predicted by the model. The parameters $\alpha_{t}$ and $\bar{\alpha}_{t}$ are predefined noise scheduling terms. In this framework, $\mathbf{x}_{0}\sim p_{data}(\mathbf{x}_{0})$ denotes a sample from the true data distribution, while $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. All formulations above hold for $t=1,2,\dots,T$.
Residual-guided denoising reformulation.
Inspired by  (Yue et al., 2023), we reformulate the forward and reverse processes to explicitly incorporate the low-resolution generation result $\tilde{\mathbf{x}}_{r}$ as a structural prior for the high-resolution generation. Instead of constructing a Markov chain solely between Gaussian noise and $p_{data}(\mathbf{x}_{0})$, we aim to build a transition path from $\tilde{\mathbf{x}}_{r}$ to the high-resolution result $\mathbf{x}_{r}$. A naive formulation such as $\mathbf{x}_{T}=\bm{\epsilon}+\tilde{\mathbf{x}}_{r}$, $\mathbf{x}_{0}=\mathbf{x}_{r}$ lacks flexibility in controlling the influence of $\tilde{\mathbf{x}}_{r}$ and $\bm{\epsilon}$. To address this, we adopt a more flexible initialization as defined in Equation (1), with $\mathbf{x}_{0}=\mathbf{x}_{r}$. Under this formulation, the residual-guided forward and reverse diffusion processes become:
@@
Equations (4) and (5) define the forward and reverse processes, respectively. The model is trained to predict $\tilde{\bm{\epsilon}}_{\theta}(\mathbf{x}_{t},t)\approx\alpha\cdot\bm{\epsilon}+\beta\cdot\tilde{\mathbf{x}}_{r}$, such that the noise component in $\mathbf{x}_{t}$ is replaced by a composition of $\bm{\epsilon}$ and the low-resolution result $\tilde{\mathbf{x}}_{r}$. Except for this reformulated initialization and noise target, the rest of the denoising process remains consistent with the original DDPM. All formulations above hold for $t=1,2,\dots,T$.
@@ 3.4.2 Controlling noise-structure balance
In the high-resolution stage, generation begins from the latent $\mathbf{x}_{T}=\alpha\cdot\bm{\epsilon}+\beta\cdot\tilde{\mathbf{x}}_{r}$, which combines a stochastic noise term and a structural prior. Importantly, although the noise term has been modified, the actual source of randomness remains from $\bm{\epsilon}$. During the reverse denoising process, this component $\alpha\cdot\bm{\epsilon}$ is gradually transformed into the residual $\mathbf{x}_{r}-\beta\cdot\tilde{\mathbf{x}}_{r}$, ensuring that the final output satisfies $\mathbf{x}_{0}=\mathbf{x}_{r}$. This formulation offers explicit control over the trade-off between randomness and structural guidance. A larger $\beta$ may cause the generation to overly rely on the low-resolution input, which is intended to provide only structural guidance, thereby potentially suppressing the recovery of fine details. While a larger $\alpha$ introduces stronger randomness that may override structural consistency. Through empirical studies, we find that setting $\alpha=\beta=0.5$ offers a good trade-off between fidelity and flexibility. More discussions can be found in Section 4.3.
@@ 4.1 Experimental setup
In this paper, we conduct experiments on the VITON-HD (Choi et al., 2021) and DressCode (Morelli et al., 2022) datasets. All ablation studies are carried out on VITON-HD. While VITON-HD contains only upper-body garments, DressCode includes three garment categories: upper-body, lower-body, and dresses. Both datasets consist of paired images, each containing a person image and a corresponding garment image. However, as our method is mask-free, the original paired data alone is insufficient for training. To address this, we use IDM-VTON (Choi et al., 2024) to synthesize additional training data. Further details are provided in Appendix A.1.
@@
For network initialization, both the reference U-Net and the denoising U-Net are initialized with pretrained weights from Stable Diffusion 1.5 (Rombach et al., 2022). As detailed in Subsections 3.3 and 3.4, we set the downsampling ratio $\sigma=2$ and use $\alpha=\beta=0.5$ to initialize $\mathbf{x}_{T}$ in the high-resolution stage. Consequently, the low-resolution stage operates at a resolution of $384\times 512$, while the high-resolution stage produces outputs at $768\times 1024$. During inference, the two stages are executed sequentially with 20 sampling steps each, using the DDIM sampler (Song et al., 2020). Additional training details are provided in Appendix A.2.
@@
Previous virtual try-on methods typically evaluate performance under both paired and unpaired settings. The paired setting involves reconstructing the original person image with the same garment, while the unpaired setting replaces it with a different one (Choi et al., 2021). As most prior approaches rely on masking the garment region, they support both settings. In contrast, our method is mask-free and is therefore evaluated only under the unpaired setting, which better reflects real-world scenarios. Following prior works (Chong et al., 2025; Jiang et al., 2024; Zhou et al., 2025), we adopt Fréchet Inception Distance (FID) (Parmar et al., 2022) and Kernel Inception Distance (KID) (Bińkowski et al., 2018) as quantitative metrics. We also conduct a user study to assess perceptual quality: for VITON-HD, we randomly sample 100 results per method; for DressCode, we sample 33, 33, and 34 results from the dresses, lower-body, and upper-body categories, respectively. Participants are asked to select the result they think performs better in the try-on task. All evaluations are conducted at a resolution of $768\times 1024$.
@@
Table 2: Ablation on dual-scale design and downsampling ratio $\sigma$.
@@
Table 3: Ablation on coefficients $\alpha,\beta$ under fixed $\sigma=2$.
@@
4.2 Quantitative and qualitative results
Qualitative comparison.
Figure 3 presents a qualitative comparison between DS-VTON and recent baseline methods on the VITON-HD (Choi et al., 2021) dataset. Row numbers referenced below correspond to Figure 3. We evaluate each method in terms of two key aspects: structural alignment and detail preservation. In terms of structural alignment, most mask-based methods fail to accurately capture body pose and garment shape, as illustrated in Row 1. FitDiT (Jiang et al., 2024), which uses a larger rectangular mask, performs relatively better but introduces noticeable artifacts: it fails to reconstruct the hands and exhibits visible artifacts at the junction of the upper and lower garments in Row 1. OOTDiffusion (Xu et al., 2025) alters the original skin tone in Row 2. CatVTON (Chong et al., 2025), IDM-VTON (Choi et al., 2024), and Leffa (Zhou et al., 2025) also show varying degrees of misalignment. In contrast, DS-VTON consistently achieves accurate alignment across a wide range of poses and garment types. Regarding detail preservation, CatVTON and IDM-VTON fail to retain fine-grained textures on complex garments (Row 3), with IDM-VTON generating oversimplified or missing patterns. While Leffa, OOTDiffusion, and FitDiT better preserve textures, they each show limitations: FitDiT achieves texture preservation at the cost of distorting the person’s actual body shape and dressing structure (Rows 2 and 4), OOTDiffusion introduces artifact patterns (Row 5), and Leffa exhibits reduced pattern clarity in complex regions (Rows 3). Furthermore, both FitDiT and Leffa suffer from tonal inconsistencies—FitDiT produces noticeably brighter garments (Row 6), while Leffa tends to generate darker outputs. In contrast, DS-VTON preserves fine textures while maintaining tonal fidelity throughout.
@@ 4.3 Ablation study
Figure 4: Visualized results under varying downsampling ratios $\sigma$. For each $\sigma$, the LR column shows the output from the low-resolution stage, and the corresponding HR column shows the high-resolution result refined from that output.
@@
To validate the effectiveness of our dual-scale design, we train four variants on the VITON-HD dataset. As shown in Table 3, Row 1 ($\sigma=1$) corresponds to training the model directly at high resolution ($768\times 1024$) without a refinement stage, aligning with the $\sigma=1$ (LR) column in Figure 5. Applying the mask-free strategy at this resolution leads to poor structural results, highlighting the need for coarse-level guidance. Row 2 ($\sigma=1,\alpha=\beta=\frac{1}{2}$), shown in the $\sigma=1$ (HR) column, adds a refinement stage. While some structural errors are alleviated, relying solely on the second stage to recover both structure and detail still results in failures, due to the lack of reliable low-resolution guidance. This also violates our design principle of decoupling structure modeling (stage one) from detail refinement (stage two). Rows 3 and 4 evaluate the full dual-scale pipeline with $\sigma=4$ and $\sigma=2$, respectively. When garments are structurally simple, both perform reasonably well. However, as shown in Row 1 of Figure 5, $\sigma=4$ introduces visible information loss in complex cases. These qualitative results align with the quantitative trends: among the four variants—single-stage ($\sigma=1$) and dual-stage ($\sigma=1$, $\sigma=2$, $\sigma=4$)—the single-stage model performs worst, $\sigma=4$ shows moderate improvement, and $\sigma=2$ achieves the best overall performance.
Ablation on coefficient-based initialization of $\mathbf{x}_{T}$.
When fixing the downsampling ratio to $\sigma=2$, the coefficients $\alpha$ and $\beta$ control how much the low-resolution output influences the initialization of the high-resolution stage. As shown in Figure 5, setting $\alpha$ too high—such as $\alpha=1,\beta=1$ or $\alpha=\tfrac{2}{3},\beta=\tfrac{1}{3}$—leads to structural distortions in the final result, despite the low-resolution output already providing accurate guidance. In both cases, the red stripe on the garment appears distorted, and in the $\alpha=\tfrac{2}{3},\beta=\tfrac{1}{3}$ configuration, the "GANT" text also becomes visibly warped. Among the remaining configurations, $\alpha=\tfrac{1}{3},\beta=\tfrac{2}{3}$ and $\alpha=\tfrac{1}{2},\beta=\tfrac{1}{2}$ yield generally good results. However, the lower-$\alpha$ version ($\alpha=\tfrac{1}{3}$) still shows minor distortion in the red stripe, indicating insufficient control over fine structure. Balancing these observations, we adopt $\alpha=\beta=\tfrac{1}{2}$ as our default setting. The quantitative results in Table 3 are consistent with this choice.
@@ 5 Conclusions
We propose DS-VTON, a dual-scale virtual try-on framework that introduces residual-guided denoising to bridge low- and high-resolution generation. This design enables a principled decoupling of the two core challenges in virtual try-on: structural alignment and detail refinement. Combined with a mask-free strategy, our approach achieves significant improvements over existing methods in both visual quality and robustness. Moreover, the proposed low-to-high resolution paradigm is inherently scalable and generalizable, offering potential for extension to higher resolutions and broader generation tasks beyond virtual try-on, such as personalized image synthesis.
Appendix
Appendix A Experimental details
A.1 Additional training data generation and associated challenges
@@
A.2 More implementation details
All experiments are conducted on 8 NVIDIA A6000 GPUs. For both VITON-HD [Choi et al., 2021] and DressCode [Morelli et al., 2022], the low-resolution stage is trained with a batch size of 8, and the high-resolution stage with a batch size of 2. Both stages are optimized using the AdamW optimizer [Loshchilov et al., 2017] with a learning rate of $1\text{e}{-6}$. For VITON-HD, the low-resolution and high-resolution stages are trained for 15,000 and 30,000 steps, respectively (approximately 5 and 24 hours). For DressCode, we jointly train all three garment categories, with the two stages trained for 20,000 and 30,000 steps (approximately 7 and 24 hours). During inference, both stages use 20 DDIM sampling steps. On a single A6000 GPU, the low-resolution stage takes about 1 second per sample, and the high-resolution stage about 4 seconds.
Appendix B More ablation studies
B.1 Ablation on removing cross-attention layers in U-Net

### v2

@@ 3.4 High-resolution stage
In this stage, $\mathbf{x}_{g}$ and $\mathbf{x}_{p}$ are used in the same way as in the low-resolution stage: $\mathbf{x}_{g}$ is passed to the reference U-Net, while $\mathbf{x}_{p}$ is encoded and concatenated with the latent, which is then input to the denoising U-Net. The main difference from the low-resolution stage lies in the initialization of the latent and the denoising process.
3.4.1 Reformulating the denoising process with blend-refine
@@
Equation 2 and Equation 3 define the forward and reverse diffusion processes, respectively. In these equations, $\bm{\epsilon},\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, and $\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)$ denotes the noise predicted by the model. The parameters $\alpha_{t}$ and $\bar{\alpha}_{t}$ are predefined noise scheduling terms. In this framework, $\mathbf{x}_{0}\sim p_{data}(\mathbf{x}_{0})$ denotes a sample from the true data distribution, while $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. All formulations above hold for $t=1,2,\dots,T$.
Blend-refine diffusion reformulation.
Inspired by (Yue et al., 2023), we reformulate the forward and reverse processes to explicitly incorporate the low-resolution result $\tilde{\mathbf{x}}_{r}$ as a structural prior for high-resolution generation. Rather than constructing a Markov chain only between Gaussian noise and $p_{\mathrm{data}}(\mathbf{x}_{0})$, we aim to build a transition path from $\tilde{\mathbf{x}}_{r}$ to the high-resolution result $\mathbf{x}_{r}$. In the original formulation of (Yue et al., 2023), $\mathbf{x}_{T}=\bm{\epsilon}+\tilde{\mathbf{x}}_{r}$ with $\mathbf{x}_{0}=\mathbf{x}_{r}$, which offers limited flexibility to control the relative influence of $\tilde{\mathbf{x}}_{r}$ and $\bm{\epsilon}$. To overcome this limitation, we introduce a simple extension with two coefficients for more flexible initialization, as defined in Equation 1, while keeping $\mathbf{x}_{0}=\mathbf{x}_{r}$. This straightforward modification yields significant improvements, as shown in Subsection 4.3. Under this formulation, the blend-refine forward and reverse diffusion processes become:
@@
Equation 4 and Equation 5 define the forward and reverse processes, respectively. The model is trained to predict $\tilde{\bm{\epsilon}}_{\theta}(\mathbf{x}_{t},t)\approx\alpha\cdot\bm{\epsilon}+\beta\cdot\tilde{\mathbf{x}}_{r}$, such that the noise component in $\mathbf{x}_{t}$ is replaced by a composition of $\bm{\epsilon}$ and the low-resolution result $\tilde{\mathbf{x}}_{r}$. Except for this reformulated initialization and noise target, the rest of the denoising process remains consistent with the original DDPM. All formulations above hold for $t=1,2,\dots,T$.
@@ 3.4.2 Controlling noise-structure balance
In the high-resolution stage, generation starts from the latent $\mathbf{x}_{T}=\alpha\cdot\bm{\epsilon}+\beta\cdot\tilde{\mathbf{x}}_{r}$, a combination of stochastic noise and a structural prior. Here, $\alpha$ controls randomness and $\beta$ controls structural guidance. Intuitively, $\beta$ should reflect the similarity between the two-scale distributions, while $\alpha$ should correspond to the size of the probability space associated with the residual term $\mathbf{x}_{r}-\beta\cdot\tilde{\mathbf{x}}_{r}$. As illustrated in Figure 2, an excessively large $\beta$ forces overreliance on the low-resolution input, suppressing fine-detail recovery, whereas a large $\alpha$ introduces noise that can disrupt structure. Empirically, setting $\alpha=\beta=0.5$ provides a balanced trade-off between fidelity and flexibility. Further discussion is provided in Subsection 4.3.
@@ 4.1 Experimental setup
In this paper, we conduct experiments on the VITON-HD (Choi et al., 2021) and DressCode (Morelli et al., 2022) datasets. All ablation studies are carried out on VITON-HD. While VITON-HD contains only upper-body garments, DressCode includes three garment categories: upper-body, lower-body, and dresses. Both datasets consist of paired images, each containing a person image and a corresponding garment image. However, as our method is mask-free, the original paired data alone is insufficient for training. To address this, we use IDM-VTON (Choi et al., 2024) to synthesize additional training data. Further details are provided in Appendix A.1.1.
@@
For network initialization, both the reference U-Net and the denoising U-Net are initialized with pretrained weights from Stable Diffusion 1.5 (Rombach et al., 2022). As detailed in Subsections 3.3 and 3.4, we set the downsampling ratio $\sigma=2$ and use $\alpha=\beta=0.5$ to initialize $\mathbf{x}_{T}$ in the high-resolution stage. Consequently, the low-resolution stage operates at a resolution of $384\times 512$, while the high-resolution stage produces outputs at $768\times 1024$. During inference, the two stages are executed sequentially with 20 sampling steps each, using the DDIM sampler (Song et al., 2020). Additional training details are provided in Appendix A.1.2.
@@
Previous virtual try-on methods typically evaluate performance under both paired and unpaired settings. The paired setting involves reconstructing the original person image with the same garment, while the unpaired setting replaces it with a different one (Choi et al., 2021). As most prior approaches rely on masking the garment region, they support both settings. In contrast, our method is mask-free and is therefore evaluated only under the unpaired setting, which better reflects real-world scenarios. We adopt Fréchet Inception Distance (FID) (Parmar et al., 2022) and Kernel Inception Distance (KID) (Bińkowski et al., 2018) as quantitative metrics. We also conduct a user study to assess perceptual quality: for VITON-HD, we randomly sample 100 results per method; for DressCode, we sample 33, 33, and 34 results from the dresses, lower-body, and upper-body categories, respectively. Participants are asked to select the result they think performs better in the try-on task. All evaluations are conducted at a resolution of $768\times 1024$.
4.2 Quantitative and qualitative results
Qualitative comparison.
Figure 3 presents a qualitative comparison between DS-VTON and recent baseline methods on the VITON-HD (Choi et al., 2021) dataset. Row numbers referenced below correspond to Figure 3. We evaluate each method in terms of two key aspects: structural alignment and detail preservation. In terms of structural alignment, most mask-based methods fail to accurately capture body pose and garment shape, as illustrated in Row 1. FitDiT (Jiang et al., 2024), which uses a larger rectangular mask, performs relatively better but introduces noticeable artifacts: it fails to reconstruct the hands and exhibits visible artifacts at the junction of the upper and lower garments in Row 1. OOTDiffusion (Xu et al., 2025) alters the original skin tone in Row 2. CatVTON (Chong et al., 2025), IDM-VTON (Choi et al., 2024), and Leffa (Zhou et al., 2025) also show varying degrees of misalignment. In contrast, DS-VTON consistently achieves accurate alignment across a wide range of poses and garment types. Regarding detail preservation, CatVTON and IDM-VTON fail to retain fine-grained textures on complex garments (Row 3), with IDM-VTON generating oversimplified or missing patterns. While Leffa, OOTDiffusion, and FitDiT better preserve textures, they each show limitations: FitDiT achieves texture preservation at the cost of distorting the person’s actual body shape and dressing structure (Rows 2 and 4), OOTDiffusion introduces artifact patterns (Row 5), and Leffa exhibits reduced pattern clarity in complex regions (Row 3). Furthermore, both FitDiT and Leffa suffer from tonal inconsistencies—FitDiT produces noticeably brighter garments (Row 6), while Leffa tends to generate darker outputs. In contrast, DS-VTON preserves fine textures while maintaining tonal fidelity throughout.
@@ 4.2 Quantitative and qualitative results
Table 2: Ablation study on dual-scale design and downsampling ratio $\sigma$.
@@
Table 3: Ablation study on coefficients $\alpha,\beta$ under fixed $\sigma=2$.
@@ 4.3 Ablation study
Figure 4: Visualized results under varying downsampling ratios $\sigma$.
@@
Here we present several ablation studies to validate the rationale of our design. In addition, we evaluate an alternative refinement method used in SDXL; see Subsection A.3 for details.
@@
To validate the effectiveness of our dual-scale design, we train four variants on the VITON-HD dataset. As shown in Table 3, Row 1 ($\sigma=1$) corresponds to training the model directly at high resolution ($768\times 1024$) without a refinement stage, aligning with the $\sigma=1$ (LR) column in Figure 4. Applying the mask-free strategy at this resolution leads to poor structural results, highlighting the need for coarse-level guidance. Row 2 ($\sigma=1,\alpha=\beta=\frac{1}{2}$), shown in the $\sigma=1$ (HR) column, adds a refinement stage. While some structural errors are alleviated, relying solely on the second stage to recover both structure and detail still results in failures, due to the lack of reliable low-resolution guidance. This also violates our design principle of decoupling structure modeling (stage one) from detail refinement (stage two). Rows 3 and 4 evaluate the full dual-scale pipeline with $\sigma=4$ and $\sigma=2$, respectively. When garments are structurally simple, both settings perform reasonably well. However, as shown in Row 1 of Figure 4, $\sigma=4$ introduces visible information loss in complex cases. These qualitative results align with the quantitative trends: among the four variants, the single stage ($\sigma=1$) and the dual stage with $\sigma=1$, $\sigma=2$, or $\sigma=4$, the single stage model performs worst, $\sigma=4$ shows moderate improvement, and $\sigma=2$ achieves the best overall performance.
Ablation on different initializations of $\mathbf{x}_{T}$.
When fixing the downsampling ratio to $\sigma=2$, the coefficients $\alpha$ and $\beta$ determine how strongly the low-resolution output influences the high-resolution stage. As shown in Figure 5, setting $\alpha$ too high, e.g., $\alpha=1,\beta=1$ or $\alpha=\tfrac{2}{3},\beta=\tfrac{1}{3}$, causes structural distortions in the final result even though the low-resolution output already provides accurate guidance. In both cases the red stripe on the garment is distorted, and in the $\alpha=\tfrac{2}{3},\beta=\tfrac{1}{3}$ configuration the “GANT” text also becomes warped. We attribute this to excessive randomness in the high-resolution stage, which then attempts to re-establish structure that should have been fully resolved in the low-resolution stage. Among the remaining settings, $\alpha=\tfrac{1}{3},\beta=\tfrac{2}{3}$ and $\alpha=\tfrac{1}{2},\beta=\tfrac{1}{2}$ produce generally good results. However, the $\alpha=\tfrac{1}{3},\beta=\tfrac{2}{3}$ setting still exhibits minor distortion in the red stripe, indicating insufficient restoration of fine details. Here, the low-resolution result exerts too strong an influence: while structurally accurate, it lacks fine details, and the high $\beta$ prevents the high-resolution stage from effectively correcting those errors. Based on these observations, we adopt $\alpha=\beta=\tfrac{1}{2}$ as our default setting. The quantitative results in Table 3 support this choice.
@@ 5 Conclusions
We propose DS-VTON, a dual-scale framework that employs blend-refine denoising to bridge low- and high-resolution generation. This design enables a more effective coarse-to-fine process and, combined with a mask-free strategy, achieves significant improvements in both visual quality and robustness over existing methods. Furthermore, the paradigm is inherently scalable and generalizable, with clear potential for extension to higher resolutions and to broader generation tasks beyond virtual try-on, such as personalized image synthesis.
Appendix A Appendix
A.1 Experimental details
A.1.1 Additional training data generation and associated challenges
@@
A.1.2 More implementation details
All experiments are conducted on 8 NVIDIA A6000 GPUs. For both VITON-HD [Choi et al., 2021] and DressCode [Morelli et al., 2022], the low-resolution stage is trained with a batch size of 8, and the high-resolution stage with a batch size of 2. Both stages are optimized using the AdamW optimizer [Loshchilov et al., 2017] with a learning rate of $1\text{e}{-6}$. For VITON-HD, the low-resolution and high-resolution stages are trained for 15,000 and 30,000 steps, respectively (approximately 5 and 24 hours). For DressCode, we jointly train all three garment categories, with the two stages trained for 20,000 and 30,000 steps (approximately 7 and 24 hours). During inference, both stages use 20 DDIM sampling steps. On a single A6000 GPU, the low-resolution stage takes about 1 second per sample, while the high-resolution stage takes about 4 seconds. Although our method includes two stages, the low-resolution stage is much faster than the high-resolution stage, so the total runtime remains acceptable. Compared with the methods in our earlier comparison, our approach is only slower than CatVTON, comparable to FitDiT, and faster than the others.
A.1.3 Additional Comparisons with Other Methods
We detail here the data requirements of the methods introduced in Subsection 4.1 to clarify the comparisons. For completeness, we also present paired-setting results on VITON-HD. All competing methods except ours require an agnostic mask; Leffa additionally uses a DensePose map; FitDiT requires human body keypoints; and IDM-VTON depends on both detailed garment descriptions and a DensePose map. Our method, by contrast, is trained only on paired images of the same person wearing different garments.
Because our approach does not mask the original garment region, the input and expected output under the paired setting are identical. Nevertheless, we include two evaluation variants:
  • DS-VTON (direct): directly uses the garment and person images to generate the result.
  • DS-VTON (train-way): first employs IDM-VTON to create an image of the person wearing another garment, then applies our method to re-dress the original garment. This procedure involves two rounds of unpaired try-on and inevitably introduces additional artifacts, so it does not fully reflect our method’s capability.
The quantitative results appear in Table 4.
A.2 More ablation studies
A.2.1 Ablation on removing cross-attention layers in U-Net